OpenStudy (anonymous):

. mc055-1.jpg Find the area of the square.

OpenStudy (anonymous):

There is no readable diagram.

OpenStudy (anonymous):

where is the square

OpenStudy (anonymous):

Click the link

OpenStudy (anonymous):

what link

OpenStudy (anonymous):

the attackment

OpenStudy (anonymous):

attachment*

OpenStudy (anonymous):

its not up there

OpenStudy (anonymous):

I just sent it again

OpenStudy (anonymous):

Let x be the square's side length. \[2 x^2 = 12^2 \]\[x^2=72 \]

OpenStudy (anonymous):

72 square root 2 yd2 72 yd2 36 square root yd2 36 yd2

OpenStudy (anonymous):

72?

OpenStudy (anonymous):

area =72